# Policy Space Response Oracles: A Survey

### v1

@@ 3.1.2 MSSs Beyond Nash
Rectified Nash
Balduzzi et al. Balduzzi et al. (2019) reformulated the strategy exploration problem as that of enlarging what they called the gamescape, which describes the strategy space covered by the empirical game. For zero-sum games, they proposed rectified Nash as an MSS designed to expand the gamescape and enhance a diversity measure called effective diversity. In rectified Nash, best responses are only applied to opponent’s equilibrium strategies that the learning player defeats or ties with.
@@
One notable observation for PSRO with NE is that the full-game regret, used as a measure for evaluating the performance of PSRO, of the empirical-game NE does not decrease monotonically over PSRO iterations. In the worst case, the full-game regrets will increase until the last iteration, when a full-game NE is found (one example can be found in the work by McAleer et al. McAleer et al. (2022b)). To address this issue, it was proposed to use minimum-regret constrained profiles (MRCP) Jordan et al. (2010); Wang et al. (2022) as an MSS. An MRCP is the profile with minimum regret with respect to the full game  [^2] MRCP was called the least-exploitable restricted distribution in the two-player zero-sum context by McAleer et al. McAleer et al. (2022b).. With MRCP as the MSS, the resulting PSRO variant is known as anytime PSRO because regret monotonically decreases as the empirical game grows. Despite the difficulty of computing MRCP in general games, anytime PSRO leverages the properties of two-player zero-sum games and computes MRCP by regret minimization against a best response (RM-BR) Johanson et al. (2012). In a further work McAleer et al. (2022a), anytime PSRO was extended by including not one but two strategies in the empirical game at each iteration, the first a best best response to MRCP and the other a best response to the other player’s latest strategy (i.e., the strategy added at the last PSRO iteration). This modification was observed to improve the performance of anytime PSRO. However, Wang et al. Wang et al. (2022) pointed out that MRCP regret will monotonically decrease for any MSS, which suggests that the changes to the MSS in anytime PSRO were not justified purely by the desire to monotonically decrease MRCP regret. We discuss the evaluation of strategy exploration further in Section 6.
@@
Apart from the issues we discussed above, Marris et al. Marris et al. (2021) further argued that NE may not be an appropriate MSS for PSRO or even a solution concept in general-sum games due to its computational intractability. Therefore, they proposed utilizing correlated equilibrium (CE) and coarse correlated equilibrium (CCE) as MSSs, introducing a variant called Joint PSRO (JPSRO). Since multiple (C)CE exist in an empirical game, they select the unique (C)CE that maximizes the Gini impurity. Theoretical analysis demonstrated that JPSRO converges to a (C)CE. Relatedly, Team-PSRO finds TMECor equilibrium in two-team zero-sum games McAleer et al. (2023b).
MSSs Motivated by Specific Games.
In addition to the above well-established solution concepts, the literature on PSRO has also investigated other solution concepts, which originate from specific games but can be generally applied to various game settings. One example is the Risk-Averse Equilibrium (RAE) introduced by Slumbers et al. Slumbers et al. (2023), aimed at managing risk in multiagent systems. Specifically, RAE minimizes potential variance in rewards by accounting for the strategies of other players. Besides, Li et al. Li et al. (2023b) presented the Nash Bargaining Solution (NBS), a concept originating from the bargaining game settings, as an MSS. NBS can be computed by maximizing the product of player utilities in the empirical game.
@@
Distinct from prior works that design MSSs based on various solution concepts and heuristics, Feng et al. Feng et al. (2021) proposed Neural Auto-Curricula (NAC) based on meta-learning, which automates the design of MSSs in an end-to-end manner. Specifically, MSSs in NAC are parametrized by a neural network, which is trained by minimizing the regret of the meta-strategy in the resulting empirical games. The resulting empirical games are generated through PSRO with the current MSS (i.e., the current neural network) in games sampled from a game distribution. With this training scheme, Feng et al. Feng et al. (2021) showed that NAC can learn an effective MSS for a family of games.
@@ 3.2 PSRO with Alternative Game Forms
Instead of employing the normal-form representation, some PSRO variants use an extensive-form representation for the empirical game, offering a richer way to encompass temporal patterns in actions and information for underlying sequential games. One such example is given by extensive-form DO (XDO) McAleer et al. (2021), a DO variant with extensive-form game models. The extensive-form empirical game tree in XDO is constructed using the restricted set of players’ strategies. Similar to DO, NE is deployed as an MSS target, computed through Counterfactual Regret Minimization Zinkevich et al. (2007), and the best response computation will result in new actions at information states in the empirical game tree. Note that when a new action is added to an information state, multiple strategies will be added. So one iteration in XDO implicitly needs more simulation for profile evaluation than one DO iteration. Periodic DO Tang et al. (2023) extends XDO by improving the stopping threshold of the restricted game solver.
In a work by Konicki et al. Konicki et al. (2022), the benefits of leveraging the extensive-form representation for the empirical game were further explored. They showed that with an extensive-form representation in PSRO, the true game can be approximated more accurately than using a normal-form model constructed from the same amount of simulation data. This accuracy improvement stems from the fact that the simulation data for modeling a chance node in extensive form can be reused while the modeling needs to be re-simulated every time for evaluating a profile in normal form.
@@
Muller et al. Muller et al. (2022) and Wang and Wellman Wang and Wellman (2023a) adapted PSRO to mean-field games (MFGs). Since the utility function for MFGs is not generally linear in the mean field, the empirical MFG model cannot be represented explicitly. Instead, Wang and Wellman Wang and Wellman (2023a) employed a game model learning approach Sokota et al. (2019), which is essentially a form of regression that learns a utility function over a restricted set of strategies and mean fields derived by these strategies. They proved the existence of NE in the empirical game with a restricted strategy set and the convergence of PSRO to NE in MFGs.
Diversity Measure Concept Diversity Type Enlargement Target Compatible MSS Behavioral Response Gamescape Policy Hull Nash $\alpha$-Rank Effective Diversity Balduzzi et al. (2019) Rectified Nash strategy Expected Cardinality Perez-Nieves et al. (2021) Determinantal point processes Convex Hull Enlargement Liu et al. (2021) Euclidean projection Occupancy Measure Mismatching Liu et al. (2021) f-divergence, occupancy measure Unified Diversity Measure Liu et al. (2022b) Strategy feature, diversity kernel Policy Space Diversity Yao et al. (2023) Bregman divergence, sequence-form
@@ 4 Strategy Exploration via RO
Besides setting MSSs, establishing ROs to guide strategy exploration in PSRO has also been explored in the literature. One predominant way to design novel ROs is to include diversity measures in the standard RO, aiming at increasing the diversity of strategies in the empirical game.
For example, Perez-Nieves et al. Perez-Nieves et al. (2021) proposed diverse PSRO, which incorporates expected cardinality, a diversity measure defined through a determinantal point process, into the standard RO. Liu et al. Liu et al. (2021) define behavioral diversity (BD) and response diversity (RD), measuring diversity on different scales. BD is defined based on the distance between action-state coverages given by different strategies, while RD measures the distance between the payoff vector induced by the new strategy and the current empirical game. Subsequently, Liu et al. Liu et al. (2022b) proposed unified diversity measures to capture a variety of diversity metrics, which were later combined with the standard RO. Yao et al. Yao et al. (2023) observed that the current diversity measures for enlarging the gamescape fail to connect to the quality of NE approximation. They connected the RO to the quality of NE approximation by introducing population exploitability (PE), which is essentially the regret of MRCP Wang et al. (2022), to reflect the coverage of a policy hull (i.e., the convex combinations of strategies in the empirical game). They showed that a larger policy hull indicates lower PE. Therefore, they proposed an RO variant to enlarge the policy hull, aiming at promoting strategy exploration. We list the specifications of different methods for enhancing diversity in Table 1.
Aside from diversity, Li et al. Li et al. (2023b) deployed Monte Carlo tree search (MCTS) as the best-response oracle using different values (e.g., social welfare) to update values of nodes along the sample path in the back-propagation step of MCTS. The employment of different back-propagation values can also be viewed as modifications of the RO.
@@ 5 Strategy Exploration via Joint MSS-RO
Generally speaking, MSSs and ROs are not functionally interchangeable and can bring a coupling impact to strategy exploration, especially when a goal is specified. In this section, we discuss prior works that jointly vary MSSs and ROs.
@@
Muller et al. Muller et al. (2020) proposed the adoption of $\alpha$-Rank Omidshafiei et al. (2019) as the preferred solution concept due to its computational scalability and uniqueness in many-player general-sum games. To make PSRO with $\alpha$-Rank as MSS converge, they introduced the preference-based best-response oracle, which essentially returns a set of strategies that maximizes the probability mass under $\alpha$-Rank from the set of better responses to the current strategy in a profile. With this MSS-RO combination, they proved that PSRO with $\alpha$-Rank converge to a concept, called sink strongly-connected components, which describes the distribution of strategies in long-term interactions.
@@
An empirical study by Wang and Wellman Wang and Wellman (2024) explicitly investigated the coupling impact of MSSs and ROs in strategy exploration. This research experimented with a multitude of MSS and RO combinations with each RO possessing unique characteristics. Their experimental results underscore the pivotal role of ROs in steering strategy exploration towards desired objectives, such as higher social welfare. Moreover, they showed that with a careful selection of MSS, the performance of strategy exploration can be further improved.
@@ 6 Evaluating Strategy Exploration
In addition to designing novel strategy exploration algorithms, a significant effort has also been put into investigating methodological considerations in evaluating strategy exploration, and proposing and justifying new evaluation methods. In PSRO, it may seem natural to employ the same MSS for both strategy generation and evaluation, as much of the prior work in PSRO exploration has done. For example, if NE is used as the MSS for strategy generation, then the regret of NE of intermediate empirical games will be used as the performance measure at each iteration of PSRO. Similarly, the regret of a uniform distribution over strategies will be the performance measure when the uniform is employed as MSS, in which case PSRO reduces to fictitious play.
Wang and Wellman Wang et al. (2022) argued this evaluation approach could yield a misleading conclusion on the performance of MSS-RO combinations. They highlighted that each MSS-RO combination essentially generates a distinct sequence of strategies, and thus the empirical game model at any point reflects a distinct strategy space. The comparisons of different MSS-RO combinations are across different strategy spaces, which may not be faithfully represented by a simple summary such as an interim solution. Therefore, they proposed to use MRCP, the profile closest to the full-game NE (in regret) in the empirical game, as the evaluation metric for evaluating the performance of multiple MSS-RO combinations. This means that the MSS employed for strategy generation is independent of the MSS (e.g., MRCP) used for evaluation, which should be fixed when comparing different empirical games, regardless of the MSS with which they are generated. The requirement of fixing the evaluation MSS is named as the consistency criteria by the authors.
@@ 7 Improvements in Training Efficiency
There are two components in PSRO that are computationally demanding: best response computation (usually achieved by deep RL) and empirical game simulation. To improve the efficiency of PSRO, various methods have been developed, addressing issues related to these two aspects.
@@
Leveraging parallelization, Lanctot et al. Lanctot et al. (2017) proposed the Deep Cognitive Hierarchy (DCH) model, which creates a training hierarchy where each player trains a best response strategy (with deep RL) against the NE of the empirical game with strategies at the same level or below it. This warm-starts best response training and speeds up PSRO compared to training best responses from scratch. Motivated by DCH, McAleer et al. McAleer et al. (2020) proposed Pipeline PSRO (P2SRO). Similar to DCH, P2SRO initializes a bunch of strategies and assigns each strategy a level. Then P2SRO warm-starts training each strategy in parallel against the NE of the empirical game involving strategies with lower levels, which accelerates the overall training of PSRO. A major contribution of P2SRO is that it optimizes the training hierarchy, addressing issues with DCH such as heavy parallelization requirements and the need for an estimate of the number of PSRO iterations.
@@
A distinctive characteristic of empirical game models is that they are derived or estimated from simulation data. To improve the sample efficiency of PSRO, Smith and Wellman Smith and Wellman (2023) proposed to learn a full-game model about the game dynamics from the simulator concurrently with running PSRO, aiming at reducing the simulation cost by querying the full-game model. Besides, Zhou et al. Zhou et al. (2022) developed an efficient PSRO (EPSRO) implementation for reducing the simulation cost of PSRO in two-player zero-sum games. The key insight is that the simulation for the empirical game is only used for computing best response target profiles. So as long as best response target profiles can be computed in other ways (e.g., a uniform MSS does not need the evaluation of a game model), there is no need to maintain the complete empirical game model, avoiding unnecessary simulations.
@@
Transfer learning is a machine learning technique where a model trained on one task is repurposed for a different task. In PSRO, best-responding to different strategies can be viewed as such tasks and thus transfer learning can be applied to warm-start training new best responses. One example leveraging this idea is NeuPL Liu et al. (2022a), which represents all strategies in the strategy set via a shared neural network. NeuPL utilizes explicit parameter sharing for skill transfer, which was shown to effectively accelerate the adaptation to the opponent’s meta-strategy. Smith et al. Smith et al. (2023) also utilized transfer learning to reduce simulation costs in computing best responses. Their Mixed-Oracle method constructs a new best response by learning and maintaining best responses to the pure strategies of the opponent (represented as Q-value functions) and then mixing (Q-values) according to the meta-strategy.

### v2

@@ 3.1.2 MSSs Beyond Nash
Rectified Nash.
Balduzzi et al. Balduzzi et al. (2019) reformulated the strategy exploration problem as that of enlarging what they called the gamescape, which describes the payoff space covered by the restricted game. For symmetric two-player zero-sum games, they proposed rectified Nash as an MSS designed to expand the gamescape and enhance a diversity measure called effective diversity. In rectified Nash, best responses are only applied to opponent’s equilibrium strategies that the learning player defeats or ties with.
@@
One notable observation for PSRO with NE is that the full-game regret of the restricted-game NE, used as a measure for evaluating the performance of PSRO, does not decrease monotonically over PSRO iterations. In the worst case, the full-game regrets will increase until the last iteration, when a full-game NE is found (one example can be found in the work of McAleer et al. McAleer et al. (2022b)). To address this issue, it was proposed to use minimum-regret constrained profiles (MRCP) Jordan et al. (2010); Wang et al. (2022) as an MSS. An MRCP is the profile with minimum regret with respect to the full game [^3] MRCP was called the least-exploitable restricted distribution in the two-player zero-sum context by McAleer et al. McAleer et al. (2022b).. With MRCP as the MSS, the resulting PSRO variant is known as anytime PSRO because regret monotonically decreases as the restricted game grows. Despite the difficulty of computing MRCP in general games, anytime PSRO leverages the properties of two-player zero-sum games and computes MRCP by regret minimization against a best response (RM-BR) Johanson et al. (2012). In a further work McAleer et al. (2022a), anytime PSRO was extended by including not one but two strategies in the restricted game at each iteration, the first a best response to MRCP and the other a best response to the other player’s latest strategy (i.e., the strategy added at the last PSRO iteration). This modification was observed to improve the performance of anytime PSRO. However, Wang et al. Wang et al. (2022) pointed out that MRCP regret will monotonically decrease for any MSS since the concept of MRCP is well-defined in any restricted game, which suggests that using MRCP as an MSS in anytime PSRO is not justified purely by the desire to monotonically decrease MRCP regret. We discuss the evaluation of strategy exploration further in Section 6.
@@
Apart from the issues discussed above, Marris et al. Marris et al. (2021) further argued that NE may not be an appropriate MSS for PSRO or even a solution concept in general-sum games due to its computational intractability. Therefore, they proposed utilizing correlated equilibrium (CE) and coarse correlated equilibrium (CCE) as MSSs, introducing a variant called Joint PSRO (JPSRO). Since multiple (C)CE exist in a restricted game, they select the unique (C)CE that maximizes the Gini impurity. Theoretical analysis demonstrated that JPSRO converges to a (C)CE. Zhao et al. Zhao et al. (2023) combined a diversity measure with (C)CE and proposed a new MSS, called diverse (coarse) correlated equilibrium (DCCE). They showed the improved performance of PSRO with DCCE over JPSRO and many other PSRO variants. Relatedly, Team-PSRO finds a team-maxmin equilibrium with coordination device Celli and Gatti (2018), an NE defined on the team level, in two-team zero-sum games McAleer et al. (2023).
Game-Motivated MSSs.
In addition to the above well-established solution concepts, the literature on PSRO has also investigated other solution concepts, which originate from specific games but can be generally applied to other game settings. One example is the Risk-Averse Equilibrium (RAE) introduced by Slumbers et al. Slumbers et al. (2023), aimed at managing risk in multiagent systems. Specifically, RAE minimizes potential variance in rewards by accounting for the strategies of other players. Another exampple is the work of Li et al. Li et al. (2023c), which proposed the Nash Bargaining Solution (NBS), a concept originating from the bargaining games, as an MSS. NBS can be computed by maximizing the product of players’ utilities in the restricted game.
@@
Distinct from prior works that design MSSs based on various solution concepts and heuristics, Feng et al. Feng et al. (2021) proposed Neural Auto-Curricula (NAC) based on meta-learning, which automates the design of MSSs in an end-to-end manner. Specifically, MSSs in NAC are parametrized by a neural network, which is trained by minimizing the regret of the meta-strategy in the resulting restricted games. These restricted games are generated through PSRO with the current MSS (i.e., the current neural network) in games sampled from a game distribution. With this training scheme, Feng et al. Feng et al. (2021) showed that NAC can learn an effective MSS for a family of games. Automated MSS design with Auto-Curricula was also discussed by Yang et al. Yang et al. (2021), who highlighted the significance of including behavioral diversity in auto-curricula and presenting several challenges in designing such auto-curricula for successful real-world applications. Another way to achieve automated MSS design is to select among existing MSSs to fit various games or different phases within a game adaptively. For example, Li et al. Li et al. (2024b) applied hyperparameter optimization to learn weights among multiple MSSs and mixed the outputs of these MSSs based on the weights as the next best response target.
@@ 3.2 PSRO with Alternative Game Forms
Instead of employing the normal-form representation, some PSRO variants use an extensive-form representation for the restricted game, offering a richer way to encompass temporal patterns in actions and information for underlying sequential games. One such example is given by extensive-form DO (XDO) McAleer et al. (2021). The extensive-form restricted game tree in XDO again only contains a subset of players’ strategies. Similar to DO, NE is deployed as an MSS target, computed through Counterfactual Regret Minimization Zinkevich et al. (2007), and the best response computation will result in new actions at information states in the restricted game tree. Note that when a new action is added to an information state, multiple strategies will be added. So one iteration in XDO implicitly needs more simulation for profile evaluation than one DO iteration. Periodic DO Tang et al. (2023) extends XDO by improving the stopping threshold of the restricted game solver.
In a work by Konicki et al. Konicki et al. (2022), the benefits of leveraging the extensive-form representation for the restricted game were further explored. They showed that with an extensive-form representation in PSRO, the true game can be approximated more accurately than using a normal-form model constructed from the same amount of simulation data. This accuracy improvement stems from the fact that the simulation data for modeling a chance node in extensive form can be reused while the modeling needs to be re-simulated every time for evaluating a profile in normal form.
@@
Muller et al. Muller et al. (2022) and Wang and Wellman Wang and Wellman (2023a) adapted PSRO to mean-field games (MFGs). Since the utility function for MFGs is not generally linear in the mean field, the restricted MFG cannot be represented explicitly. Instead, Wang and Wellman Wang and Wellman (2023a) employed a game model learning approach Sokota et al. (2019), which is essentially a form of regression that learns a utility function over a restricted set of strategies and mean fields derived by these strategies. They proved the existence of NE in the restricted MFG and the convergence of PSRO to NE in MFGs.
Diversity Measure Concept Diversity Type Enlargement Target Compatible MSS Behavioral Response Gamescape Policy Hull Nash $\alpha$-Rank Effective Diversity Balduzzi et al. (2019) Rectified Nash strategy Expected Cardinality Perez-Nieves et al. (2021) Determinantal point processes Convex Hull Enlargement Liu et al. (2021) Euclidean projection Occupancy Measure Mismatching Liu et al. (2021) f-divergence, occupancy measure Unified Diversity Measure Liu et al. (2022c) Strategy feature, diversity kernel Policy Space Diversity Yao et al. (2023) Bregman divergence, sequence-form
@@ 4 Strategy Exploration via RO
Besides setting MSSs, establishing ROs to guide strategy exploration in PSRO has also been explored in the literature. One predominant way to design novel ROs is to include diversity measures in the standard RO, aimed at increasing the diversity of strategies in the restricted game. The importance of maintaining a diverse population of strategies has been demonstrated by many works in various domains Czarnecki et al. (2020); Zahavy et al. (2023).
Specifically, Perez-Nieves et al. Perez-Nieves et al. (2021) proposed diverse PSRO, which incorporates expected cardinality, a diversity measure defined through a determinantal point process, into the standard RO. Liu et al. Liu et al. (2021) define behavioral diversity (BD) and response diversity (RD), measuring diversity on different scales. BD is defined based on the distance between action-state coverages given by different strategies, while RD measures the distance between the payoff vector induced by the new strategy and the current restricted game. Subsequently, Liu et al. Liu et al. (2022c) proposed unified diversity measures to capture a variety of diversity metrics, which were later combined with the standard RO. Yao et al. Yao et al. (2023) observed that the current diversity measures for enlarging the gamescape fail to connect to the quality of NE approximation. They connected the RO to the quality of NE approximation by introducing population exploitability (PE), which is essentially the regret of MRCP Wang et al. (2022), to reflect the coverage of a policy hull (i.e., the convex combinations of strategies in the restricted game). They showed that a larger policy hull indicates lower PE. Therefore, they proposed an RO variant to enlarge the policy hull, aiming at promoting strategy exploration. We list the specifications of different methods for enhancing diversity in Table 1.
Aside from diversity, Li et al. Li et al. (2023c) deployed Monte Carlo tree search (MCTS) as the best-response oracle using different values (e.g., social welfare) to update values of nodes along the sample path in the back-propagation step of MCTS. The employment of different back-propagation values can also be viewed as modifications of the RO.
@@ 5 Strategy Exploration via Joint MSS-RO
Generally speaking, MSSs and ROs have a coupled impact on strategy exploration. In this section, we discuss prior works that jointly vary MSSs and ROs.
@@
Muller et al. Muller et al. (2020) proposed the adoption of $\alpha$-Rank Omidshafiei et al. (2019) as the preferred solution concept due to its computational scalability and uniqueness in many-player general-sum games. To make PSRO with $\alpha$-Rank as MSS converge, they introduced the preference-based best-response oracle, which essentially returns a set of strategies that maximizes the probability mass under $\alpha$-Rank from the set of better responses to the current strategy profile. With this MSS-RO combination, they proved that PSRO with $\alpha$-Rank converges to something that they call a sink strongly-connected component, which describes the distribution of strategies in long-term interactions. Yang et al. Yang et al. (2020) highlighted that as the number of players significantly increases, the Markov chain required in the computation of $\alpha$-Rank becomes prohibitively large, yielding a scalability problem for $\alpha$-Rank. To handle this challenge, they developed an efficient implementation of $\alpha$-Rank based on DO and stochastic optimization.
@@
An empirical study by Wang and Wellman Wang and Wellman (2024) explicitly investigated the coupling impact of MSSs and ROs in strategy exploration. This research experimented with many MSS-RO combinations with unique characteristics. Their experimental results underscore the pivotal role of ROs in steering strategy exploration towards desired objectives, such as higher social welfare. Moreover, they showed that with a careful selection of MSS, the performance of strategy exploration can be further improved.
@@ 6 Evaluating Strategy Exploration
In addition to designing novel strategy exploration algorithms, a significant effort has also been put into investigating methodological considerations in evaluating strategy exploration, and proposing and justifying new evaluation methods. In PSRO, it may seem natural to employ the same MSS for both strategy generation and evaluation, as much of the prior works in PSRO exploration have done. For example, if NE is used as the MSS for strategy generation, then the regret of NE of intermediate restricted games will be used as the performance measure at each iteration of PSRO. Similarly, the regret of a uniform distribution over strategies will be the performance measure when the uniform distribution is employed as MSS, in which case PSRO reduces to fictitious play.
Wang and Wellman Wang et al. (2022) argued that this evaluation approach could yield a misleading conclusion about the performance of MSS-RO combinations. They highlighted that each MSS-RO combination essentially generates a distinct sequence of strategies, and thus the restricted game at any point reflects a distinct strategy space. The comparisons of different MSS-RO combinations are across different strategy spaces, which may not be faithfully represented by a simple summary such as an interim solution. Therefore, they proposed to use the regret of MRCP, where MRCP is the profile closest to the full-game NE (in regret) in the restricted game, as the evaluation metric for evaluating the performance of multiple MSS-RO combinations. This means that the MSS employed for strategy generation is independent of the MSS (e.g., MRCP) used for evaluation, which should be fixed when comparing different restricted games, regardless of the MSS with which they are generated.
@@ 7 Improvements in Training Efficiency
There are two components in PSRO that can be computationally demanding: best response computation and payoff simulation in the restricted game. To improve the efficiency of PSRO, various methods have been developed, addressing issues related to these two aspects.
@@
Leveraging parallelization, Lanctot et al. Lanctot et al. (2017) proposed the Deep Cognitive Hierarchy (DCH) model, which creates a training hierarchy where each player trains a best response strategy (with deep RL) against the NE of the restricted game with strategies at the same level or below it. This warm-starts best-response training, and speeds up PSRO compared to training best responses from scratch. Motivated by DCH, McAleer et al. McAleer et al. (2020) proposed Pipeline PSRO (P2SRO). Similar to DCH, P2SRO initializes a bunch of strategies and assigns each strategy a level. Then P2SRO warm-starts training each strategy in parallel against the NE of the restricted game involving strategies with lower levels, which accelerates the overall training of PSRO.
@@
A distinctive characteristic of restricted games is that they are derived or estimated from simulation data. To improve the sample efficiency of PSRO, Smith and Wellman Smith and Wellman (2023) proposed to learn a full-game model about the game dynamics from the simulator concurrently with running PSRO, aimied at reducing the simulation cost by querying the full-game model. Zhou et al. Zhou et al. (2022) developed an efficient PSRO (EPSRO) implementation for reducing the simulation cost of PSRO in two-player zero-sum games. The key insight is that the simulation for the restricted game is only used for computing best-response target profiles. So as long as best-response target profiles can be computed in other ways (e.g., a uniform MSS does not need the evaluation of a restricted game), there is no need to maintain the complete restricted game, avoiding unnecessary simulations.
@@
Transfer learning is a machine learning technique where a model trained on one task is repurposed for a different task. In PSRO, best-responding to different strategies can be viewed as such tasks and thus transfer learning can be applied to warm-start training new best responses. One example leveraging this idea is NeuPL Liu et al. (2022b), which represents all strategies in the strategy set via a shared neural network. NeuPL utilizes explicit parameter sharing for skill transfer, which was shown to effectively accelerate the adaptation to the opponent’s meta-strategy. Liu et al. Liu et al. (2022a) further generalized NeuPL by optimizing best-responses against mixed-strategy profiles randomly sampled from the current restricted game, offering approximate optimality against any mixture over a diverse set of strategies at test time. Liu et al. Liu et al. (2024) combined NeuPL and JPSRO, enabling the transfer of knowledge in the computation of (C)CE with PSRO. Smith et al. Smith et al. (2023) also utilized transfer learning to reduce simulation costs in computing best responses. Their Mixed-Oracle method constructs a new best response by learning and maintaining best responses to the pure strategies of the opponent (represented as Q-value functions) and then mixing (Q-values) according to the meta-strategy.